OpenStudy (anonymous):

Given the function g(x) = −4x + 5, find g(−4).

OpenStudy (anonymous):

g(-4) just means g(x) when x=-4. Does this help?

OpenStudy (ajprincess):

substitute x with -4 nd add. u will get the answr.

OpenStudy (anonymous):

g(-4)=-4*(-4) + 5

OpenStudy (anonymous):

can you tell me what g(2) =?

OpenStudy (anonymous):

It would equal -3, right?

OpenStudy (anonymous):

yep =]

OpenStudy (anonymous):

So the answer is 13. Thanks!

OpenStudy (anonymous):

(-4)(-4)+5=16+5 = 21